User name jose
Log entry time 02:27:13 on December 9,2007
Entry number 219197
keyword=Hall A tools snapshot
At 02:15 both spectrometers are back on line and settled down. There are two RED readings for left D-G dipole (99.7) and rigth Q3 (100.1), presumably because of too high values.